OpenStudy (anonymous):

Consider this equation in x, y, and z: 7 = 3x + 2y - z How many more equations would you need to solve x, y, and z? In which variables can the additional equations be? Give examples of equations that would help solve these variables.

OpenStudy (anonymous):

you have 3 variables so you need 3 equations

OpenStudy (anonymous):

but what would they be, how would i figure out what the equations would be?

OpenStudy (anonymous):

the other 2 equations can anything

OpenStudy (anonymous):

ex. x=1, y=1 or x=y, y=1, or x+y=100, y+z= 1000 or x+y-z=-100000, z=1

OpenStudy (anonymous):

how would those equations help to solve for x,y and z though?

OpenStudy (anonymous):

through the use of substitution or elimination method...

OpenStudy (anonymous):

so are you creating a system of equations by putting more equations in there?

OpenStudy (anonymous):

short answer, YES what you're given is a single equation a system of equation refers to more than 1 equation that could mean 2 equations or 3 equations or 100 equations

OpenStudy (anonymous):

they're sort of classified by the number of equations and the number of variables ex. "2 equations 2 unknowns" "5 equations 4 unknowns" general rule of thumb, if the number of variables is more than the number of equations then you cannot completely solve the system of equations if the number of variables are less than the number of equations, then a. something is wrong b. one of the equations are repeated ex. x+y=1 is the same as 2x+2y=2

OpenStudy (anonymous):

to completely solve for a system of equations, you need at least "n equations n unknowns" the number of equations must be equal to the number of variables note the word AT LEAST